OpenStudy (anonymous):

Can somebody help me with 2 algebra questions please? 1. Add and answer in simplest fractional form. -1/5 + -3/4 = (Points : 1) a. -4/9 b. -19/20 c. -4/20 2.Add and answer in simplest fractional form. -4/3 + 7/8 = (Points : 1) a. -11/24 b. 3/5 c. 3/24

OpenStudy (anonymous):

the answer for number one can be b or c and the aswer for number 2 can be a or b

OpenStudy (anonymous):

first get the least common denominator between the two and then add. 1) -1/5 = -4/20 and -3/4=-15/20 so add -4/20+(-15/20)=-19/20 so the answer is B 2) -4/3=-32/24 and 7/8=21/24 so add -32/24+21/24=-11/24 so the answer is A

OpenStudy (anonymous):

@Alikat @seahawks1021 Thank you^^

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

no problem